find the partial fraction expansion of

993 +6192 + 275s + 375 s4 +993 +43s2 + 75s

pay attention, the final answer should not have complex or imaginary numbers.
Answer #1

consider 1. fcs) = 95 +615²+2755.+375 - s4 +993 +4352+755 ->F(1) = 958 +6152+27554375 5(83+9524435+75) → FCS) = 950 +618? +27Equating coefficients of powers of s on both sides, we get AtB+C = 9. 9 A+GB+ C+D=61 - (3) 43A +253 +30= 375 — 4). 75A= 375 =substitute B=3 in ean (8), we get D = 20- 25x2= 20-25 =-5. s-5 Str Now plugging the value of A, B, C and Din egna), we get f(